use the distributive property to write the expression as an equivalent expression. then evaluate it (-2-4)(-10) answers a:20+40;60 b:-2+40;38 c:(-6)(-10);-16 d:-12+14:2
Answer by nyc_function(2733)

Let me define the distributive property.
Distributive property is a(b + c) = ab + ac.
What the above law means is to multiply every item inside a parentheses or bracket by the number or letter that is "standing" outside.
However, in your question, there is no need to apply the distributive property.
Inside the parentheses we have (-2-4) = (-6).
We now multiply (-6) by (-10) to get 60.
The answer is choice a.
If you want to use the law, then we do it this way:
(-2-4)(-10) = (-2)(-10) + (-4)(-10) = 20 + 40 = 60.
Understand?